A company will manufacture a total of 10,000 units of its products at plants A and B. Available data are as follows:

Plant A Plant B
Variable cost per item $6 $6.50
Fixed cost $40,000 $35,000

The company decided to spend no more than $137,500 for total costs. What is the maximum number of units that must be produced at Plant B?
Answer #1

Total units = 10,000

Let, Units produced for Plant A = a

then, Units produced for plant B = 10000-a

Cost for plant A = $40,000 + $6*a

Cost for plant B = $35,000 + $6.50*(10,000-a)

40,000+6a+35,000+65000-6.5a = 137500

140000-0.5a = 137500

a = 5,000 units

Maximum number of units that must be produced at Plant B = 10000-5000 = 5000 units